OpenStudy (anonymous):

for f(x)=2x^2+2x-6, find f(k-1)

OpenStudy (anonymous):

put x=k-1

OpenStudy (lgbasallote):

just replace x with k-1 \[\implies f(k-1) = 2(k-1)^2 + 2(k-1) - 6\] simplify to get the final answer does that help?

OpenStudy (anonymous):

Yes that does! I wasnt sure if that was all I had to do

OpenStudy (lgbasallote):

like i said..you still ahve to expand it